# Synthetic Data Generation for Culturally Nuanced Commonsense Reasoning in Low-Resource Languages

### v1

@@ 4.1 Overall Statistics
For LLM-assisted dataset creation, GPT-4o and Claude model demonstrated the highest efficiency, generating nearly 1,000 clean samples with minimal discarded output, while Mixtral was the least efficient, requiring significantly more samples to reach the same threshold (see Appendix F). The LLM-generated data does not tend to have a uniform topic distribution due to variations in broken samples.
In total, the LLM-generated datasets contain 72K sentences and approximately 557K words. The word distribution across sentence positions remains consistent across the six LLMs, with word counts per position relatively uniform and a median sentence length ranging from 5 to 10 words (see Appendix E).
The MT and Human-written datasets exhibit a similar word distribution pattern to the LLM-generated datasets. MT contains around 6K sentences with 44,5K words, while Human-written data has 6,7K sentences with 58,2K words. Despite the slight difference in word count, both datasets maintain a consistent distribution, with a median sentence length ranging from 4 to 11 words.
Model Sundanese Javanese
@@
GPT-4o 4.7 4.2 80 96 4.9 4.5 96 68
Claude 4.7 4.4 86 92 4.9 4.3 92 74
Cohere 3.4 3.0 28 46 4.5 4.1 70 52
Llama3.1 3.7 3.4 56 70 4.5 4.3 54 44
Gemma2 3.9 3.1 42 78 4.8 3.6 76 54
Mixtral 2.0 1.5 0 4 2.2 2.1 0 6
MTtest 5.0 5.0 100 14 5.0 5.0 100 18
HW 5.0 5.0 100 96 5.0 5.0 100 66
Table 2: Quality analysis of various models on Sundanese and Javanese. Higher scores indicate better performance in each category. “HW” denotes human-written dataset
Model Sundanese Javanese
@@
MTtest 500 3877 0 0.82 500 3620 3 0.81
HW 594 4693 0 0.84 529 3497 0 0.80
Table 3: Lexical diversity analysis of different models on Sundanese and Javanese.
@@ 4.2 Quality Analysis based on Human Evaluation
We evaluate the quality of LLM-generated, machine-translated, and human-written datasets through human assessment based on four key criteria: coherence, fluency, correctness, and cultural relevance. Specifically, we randomly select 50 samples (5–10% of the total dataset) for both Sundanese and Javanese and engage two native speakers for evaluation. Coherence and fluency are rated on a Likert scale from 0 to 5, while correctness and cultural relevance are assessed using binary annotation and reported as percentages. More details on the annotation guideline can be found in Appendix J. Table 2 summarizes the results of human evaluation, with scores for coherence and fluency being averaged between the annotators, while we count the percentage of data perceived as correct or culturally relevant by both annotators to aggregate scores on correctness and cultural relevance. The Pearson correlation between annotators for coherence and fluency ranges from 0.6 to 0.7, while Kappa scores for correctness and cultural relevance range from 0.4 to 0.6.
We observe that across LLMs, GPT-4o and Claude model consistently demonstated strong performance across all metrics, especially in correctness and cultural relevance. Notably, in terms of cultural relevance, both models achieved cultural relevance scores comparable to human written text, making them the best-performing LLMs in this aspect. This also indicates that GPT-4o and Claude model are capable of generating culturally appropriate content at a level similar to human writing. However the major issue is in correctness, where GPT-4o has 20% error in Sundanese. Machine translation with post-edit might produce a good coherence, fluency, and correctness. However, the data has lower cultural relevance score. This highlight that traditional approach with machine translation is not the best practice in low-resource
@@ 4.3 Lexical Diversity
As shown in Table 3, GPT-4o achieves a MATTR score of 0.82 for Sundanese and 0.80 for Javanese, making it highly comparable to human-written data. Claude also demonstrates strong lexical diversity, maintaining a rich vocabulary while avoiding borrowed words. Machine translation (MTtest) has a competitive MATTR score but contains 3% loanwords in Javanese.
Among the LLM-generated datasets, Mixtral has the largest vocabulary size but also introduced the highest proportion of loanwords, with 16% in Sundanese and 13% in Javanese, suggesting a significant reliance on non-native terms. In contrast, GPT-4o and Claude generate text entirely in Sundanese and Javanese without incorporating foreign words, highlighting their ability to produce better datasets.
In name generation, LLMs frequently produce common Javanese names as Ayu, Dwi, Bayu, Eko, and Sari, while Mixtral and Gemma2 showed greater diversity rather than other LLM. For Sundanese, LLMs primarily generate names like Lia, Budi, Dewi, Rina, whereas Llama3.1 and Gemma2 introduce more varied name choices. All these names are widely used in Indonesia. However, variations in honorific terms are minimal across all models, with only Pak and Bu (along with variants like Ibu and Bapak) appearing consistently.
@@ 5.1 Classification
Setup
We adopt the classification accuracy metric, as proposed in both Mostafazadeh et al. (2016) and Koto et al. (2022a), defined as the ratio of correctly classified cases to the total number of test cases, $\frac{correct}{testcases}$. We fine-tune all models using the training set in Javanese and Sundanese combined, and average the results over three runs to ensure robustness. In this study, we train several models and evaluate their adaptability in the two languages. We employ several models for our experiments: BiLSTM, and XLM-R (Conneau et al., 2019) (Details in Appendix A).
Training #Data BiLSTM XLM-R
ID 1K 47.7 ± 1.1 62.8 ± 0.4
MTTrain 2K 58.2 ± 0.6 69.4 ± 0.5
GPT-4o 2K 62.8 ± 0.6 72.7 ± 0.5
Claude 2K 64.9 ± 0.6 72.0 ± 0.8
Cohere 2K 57.1 ± 0.1 68.6 ± 0.8
Llama3.1 2K 60.1 ± 0.6 72.0 ± 0.4
Mixtral 2K 60.2 ± 0.5 54.9 ± 0.5
Gemma2 2K 55.3 ± 0.9 56.6 ± 0.7
All LLMs 12K 65.1 ± 0.1 72.5 ± 0.4
All LLMs + Filtering 1.2K 62.7 ± 0.3 70.4 ± 0.2
All LLMs + Filtering 0.6K 61.6 ± 0.7 69.4 ± 0.1
Table 4: Accuracy across different training sets. “All LLMs + Filtering” has two variants: (1) a curated set of 1,220 samples selected through the automatic filtering pipeline and (2) a subset of 674 samples containing only filtered data generated by GPT-4o and Llama3.1.
Overall Performance
Table 4 shows that training on LLM-generated datasets mostly improves performance over the baseline ID dataset, highlighting the decent quality of synthetic data for both languages. Among the LLM-generated training sets, Claude and GPT-4o perform particularly well, 3–4 points higher than machine translation (MTTrain). We also observe that data generated by Mixtral and Gemma2 yields worse accuracy than MTTrain. The full results of the classification experiments are shown in Appendix I
Train with Filtered Data
In As shown in Table 4, when using the full LLM-generated dataset (12K samples), XLM-R achieves an accuracy of 72.5, which is competitive with GPT-4o. After applying filtering, the training set is reduced to 1,220 and 674 samples (61% and 33.7% of the total training data generated by GPT-4o). Despite this substantial reduction in data size, accuracy drops by only 1–3%, highlighting the efficiency of training with a smaller yet more representative subset.
@@
Accuracies on Machine-translated and Human-written Test Set
Figure 2 shows that when XLM-R is trained on MTTrain, it achieves similar accuracy (70%) on both MTTest and HW. Interestingly, compared to LLM-generated training data, MTTrain performs comparably or even better on MTTest but worse on HW. This is likely because MTTrain and MTTest originate from Koto et al. (2022b), resulting in a similar topical distribution. In contrast, the human-written test set and the LLM-generated training data were created using the same predefined topics.
Figure 3: Classification accuracy across topics.
Performance across Diferrent Topics
Figure 3 shows the performance of XLM-R in different topics. The chart shows that the data generated by GPT-4o and Claude consistently achieve high accuracy in most categories, but they also still underperform in some. The two models exhibit strong accuracy in categories such as Art, Socio-religious aspects of life, Agriculture and Wedding, and are generally weaker in categories like Fisheries and Trade, Death, Daily Activities and Pregnancy and Kids. It can also be seen that the two models give almost identical performance in Sundanese whereas GPT-4o seems slightly better in Javanese.
@@ 5.2 Generation
Setup
We use a four-sentence premise as input and train Indonesian LLM, Sahabat.AI (GoToCompany, 2024), to generate the fifth sentence in both Javanese and Sundanese. We use QLoRA (Dettmers et al., 2023) to perform supervised fine-tuning. The hyperparameters for training can be seen on Appendix A. Automatic evaluation methods are used to evaluate the model performance: ROUGE-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and BERTScore (Zhang et al., 2020).
Overall Performance
Table 5 presents the automated evaluation metrics for different training sets across Sundanese and Javanese. Training on LLM-generated datasets generally improves performance over the baseline ID dataset. However, MT achieves the highest scores in Sundanese for ROUGE-L and METEOR and outperforms all LLM-generated datasets across all metrics in Javanese. Among LLM-generated datasets, Llama3.1 performs best, particularly in Javanese, where it achieves the highest ROUGE-L and METEOR scores.
Train Set Sundanese Javanese
R-L M BS R-L M BS
ID 17.8 12.0 70.7 19.9 13.7 72.1
MT 19.1 13.4 70.4 21.8 16.7 72.5
Claude 18.1 12.7 71.6 20.4 15.5 72.2
Llama3.1 18.6 12.7 70.5 23.2 18.0 72.6
Table 5: Generation performance across different training sets for Sundanese and Javanese test sets. Metrics: ROUGE-L (R-L), METEOR (M), and BERTScore (BS).
Figure 4: Human evaluation on generation task. LLM means model trained with LLM-generated data
Human Evaluation
We conducted a human evaluation to compare models trained on machine-translated data with those trained on LLM-generated data, selecting Claude as the representative LLM model. Workers were presented with two randomized responses and asked to determine whether the first response was better, the second was better, or if both were equally good (draw). The results in Figure 4 indicate a slight preference for the model trained on LLM-generated data over MTTrain in both languages.
@@ 7 Limitations
Cultural Nuance, Language, and Dialect Coverage This study acknowledges inherent limitations in the scope of its cultural nuance, language, and dialect coverage. While we carefully curated our data and covered 12 distinct cultural topics, our data sources and predefined topics may not fully capture the comprehensive range of cultural expressions within these multicultural languages, encompassing all dialects and regional variations. Furthermore, while we performed manual human evaluation to assess the quality of our generated data to some extent, the evaluation was limited in scale and depth. Future research should explore a broader array of cultural themes, incorporate diverse data sources, improve human evaluation methods, and deepen the dialectal representation to better reflect the diversity of the targeted languages. We also suggest future studies extends the research to other local languages in Indonesia.
Prompt and Hyperparameter Variation Due to resource constraints and the primary focus on data generation strategies, this research explored a limited range of LLM prompts and hyperparameter configurations. While prompt engineering and model training were performed with due diligence, it is important to acknowledge that alternative prompts or hyperparameter settings could potentially yield different results. Future research should investigate a wider spectrum of prompts and hyperparameters to identify configurations that maximize the performance of LLMs in generating culturally nuanced common sense reasoning in Javanese and Sundanese.
@@ Appendix A Training Configurations
A.1 Classification
  • n-gram overlap: For this approach, we select the candidate with the highest overlap of n gram, measured by ROUGE-1 (Lin, 2004), calculated between the premise and the end. This metric ensures optimal alignment between the candidate and the reference text.
  • FastText-similarity: We used a similarity-based classification method, selecting the candidate with the highest cosine similarity between the premise and the ending, using the 300-dimensional FastText word embeddings trained in respective local languages (Mikolov et al., 2018). [^7] https://fasttext.cc/docs/en/crawl-vectors.html
  • Hierarchical BiLSTM: We used a two-level 200-dimensional BiLSTM architecture similar to Koto et al. (2022a), with 300-dimensional FastText representations of the respective language as input.
  • Pre-trained Language Models (PLM): We fine-tuned XLM-R (Conneau et al., 2019), a pre-trained multilingual language model.
For the LSTM model, we limited each sentence to a maximum of 30 tokens and trained it for up to 100 epochs, using early stopping with patience of 20 epochs. The model was trained with a batch size of 20, using the Adam optimizer and a learning rate of 0.1.
We set the maximum token length for the pre-trained language model to 450 for the premise and 50 for the ending sentence. The model was trained over 20 epochs with early stopping (patience set to 5), using a batch size of 40, Adam optimizer, an initial learning rate of 5e-6 for XLM-R, and a warm-up phase comprising 10% of the total training steps.
A.2 Generation
We do supervised fine-tuning on Gemma2-9B-CPT-Sahabat-AI-v1-base (GoToCompany, 2024) base model using QLoRA. The model was trained with 4 bit quantization, lora rank 64, lora alpha 128, using a batch size of 8, gradient accumulation of 8, learning rate 2e-4, with 1 epoch. We use Unsloth.ai to perform the training (Daniel Han and team, 2023).
A.3 Data Filtering Classifier
We train XLM-R model with the maximum token length of 1024 for the premise and 128 for the ending sentence. The model was trained over 26 epochs, using a batch size of 16, Adam optimizer, an initial learning rate of 1e-5 for XLM-R, and a warm-up phase comprising 5% of the total training steps.
Appendix B Prompt Template
@@
Appendix C Filtered LLM-Generated Train Data Proportion
Initially, training data was generated using six different LLMs, with each model contributing approximately 16.67% of the total 12,000 samples (around 2,000 samples per model). However, after filtering the bad examples, the final dataset composition shifted. The data consist of 1,220 samples, the distribution is as follows: Claude (37.7%), GPT-4o (29.0%), LLama (8.6%), Cohere (4.3%), Mixtral (2.7%), and Gemma-2 (1.7%).
Appendix D Topic Distribution
Figure 5 illustrates the balanced distribution of topics in various datasets generated by different LLM (Claude, Cohere, GPT-4o, Mixtral, Llama3.1 and Gemma2). Each topic contains approximately 166 samples, with minor variations, ensuring an even distribution across all categories. Meanwhile, Figure 6 visualizes the human-written dataset, where topics are similarly well distributed across both local languages.
Figure 5: Topic Distribution Across Datasets
Figure 6: Topic Distribution in Human-Written Data
Appendix E Sentence Length Distribution
Figure 8 illustrates the distribution of sentence lengths across six datasets generated by various LLMs. Each dataset exhibits a similar distribution, with most sentence lengths ranging from 5 to 10 words. While Figure 7 shows the distribution of sentence lengths from MT and Human-written, and both dataset exhibit a similar pattern of distribution as LLM-generated data, with most sentence lengths are ranging from 4 to 11.
Figure 7: Sentence Length Distribution Across Datasets: The figure shows the distribution of sentence lengths across MT and Human-written. Each bar represents the frequency of sentence lengths for different sentence types in each dataset.
Figure 8: Sentence Length Distribution Across Datasets: The figure shows the distribution of sentence lengths across various datasets generated by different LLMs. Each bar represents the frequency of sentence lengths for different sentence types in each dataset.
Appendix F Overgeneration Distribution
Figure 9 shown, the number of generated samples needed to produce 1,000 clean training data for Javanese and Sundanese varies between models, where lower numbers indicate higher efficiency due to fewer unusable examples such as duplicates or repetitions of premises.
Figure 9: Samples required to generate 1,000 clean training data: lower values indicate higher efficiency.
Appendix G Top-5 Bigrams Trigrams
Table 6 presents the top-5 bigrams and trigrams found in the dataset, highlighting frequent patterns such as "Pak Budi" and "warga desa" (villagers), reflecting common themes in LLM-generated Javanese and Sundanese texts.
Bigram (unique: 103,022) Freq (%)
pak budi (Mr. Budi) 0.2488
seneng banget (really happy) 0.2136
warga desa (villagers) 0.1660
lan kanca-kancane (and their friends) 0.1406
dina acara (event day) 0.1077
Trigram (unique: 167,659) Freq (%)
seneng banget karo (very happy with) 0.670
warga desa padha (villagers will) 0.440
dina budaya sunda (in Sundanese culture) 0.417
nu aya di (the one in) 0.335
kabeh warga desa (every villager) 0.326
Table 6: Top bigram and trigram occurrences in the dataset.
Appendix H Agreement Score
We calculated the fluency and coherency agreement scores using Pearson’s correlation and computed the correctness and cultural relevance scores using Cohen’s kappa. To evaluate the quality of LLM-generated examples, the workers followed a predefined scoring guideline (Appendix J). The agreement scores for Javanese and Sundanese are summarized in Table 7.

### v2

@@ 4.1 Overall Statistics
For LLM-assisted dataset creation, GPT-4o and Claude demonstrated the highest efficiency, generating nearly 1,000 clean samples with minimal discarded output, while Mixtral was the least efficient, requiring significantly more samples to reach the same threshold. The LLM-generated data does not have a uniform topic distribution due to variations in broken samples.
In total, the LLM-generated datasets contain 72K sentences and approximately 557K words. The word distribution across sentence positions remains consistent across the six LLMs, with word counts per position relatively uniform and a median sentence length ranging from 5 to 10 words. MTtest+Human (both with and without localization) and HW datasets exhibit a similar word distribution pattern to the LLM-generated datasets. MT contains around 6K sentences with 44,5K words, while HW has 6,7K sentences with 58,2K words. Despite the slight difference in word count, both datasets maintain a consistent distribution, with a median sentence length ranging from 4 to 11 words.
Table 1: Quality analysis of models on Sundanese and Javanese. Higher scores indicate better performance in each category.
Dataset Sundanese Javanese
@@
Human Written (HW) 5.0 5.0 100 96 5.0 5.0 100 66
LLM Generated Data
GPT-4o 4.7 4.2 80 96 4.9 4.5 97 91
Claude 4.7 4.4 86 92 4.9 4.3 96 93
Cohere 3.4 3.0 28 46 4.6 4.1 80 65
Llama3.1 3.7 3.4 56 70 4.5 4.2 65 50
Gemma2 3.9 3.1 42 78 4.8 3.6 83 81
Mixtral 2.0 1.5 0 4 2.0 1.9 3 22
Translated
MTtrain 3.24 4.36 80 20 4.36 4.46 98 12
Translated Data + Localization
MTtrain + GPT4o 4.36 4.68 86 80 4.6 4.64 98 76
MTtest + Human 5.0 5.0 100 14 5.0 5.0 100 18
Table 2: Lexical diversity analysis of different models on Sundanese and Javanese. “LW” denotes the percentage of loanword.
Dataset Sundanese Javanese
@@
Translated Data
MTtrain 1000 5272 0 0.83 1000 5007 1 0.81
Translated Data + Localization
MTtrain + GPT4o 1000 4985 1 0.82 1000 4692 1 0.81
MTtest + Human 500 3877 0 0.82 500 3620 3 0.81
@@ 4.2 Quality Analysis based on Human Evaluation
We evaluate the quality of all the constructed dataset based on four key criteria: coherence, fluency, correctness, and cultural relevance. Specifically, we randomly select 50 samples (5–10% of the total dataset) for both Sundanese and Javanese and engage two native speakers for evaluation. Coherence and fluency are rated on a Likert scale from 0 to 5, while correctness and cultural relevance are assessed using binary annotation and reported as percentages. More details on the annotation guideline can be found in Appendix D. Table 1 summarizes the results of human evaluation, with scores for coherence and fluency being averaged between the annotators. Meanwhile, for correctness and cultural relevance, we count the percentage of data perceived as correct or culturally relevant by both annotators. Inter-annotator agreement scores ranges from 0.4 to 0.7, as shown in Appendix C.
[…]
@@ 4.3 Lexical Diversity
As shown in Table 2, GPT-4o achieves a high MATTR score, making it highly comparable to human-written data. Among the LLM-generated datasets, Mixtral has the largest vocabulary size but also introduced the highest proportion of loanwords, with 16% in Sundanese and 13% in Javanese, suggesting a significant reliance on non-native terms. In contrast, GPT-4o and Claude generate text entirely in Sundanese and Javanese without incorporating foreign words, highlighting their ability to produce better datasets.
Upon manual inspection, we find that LLMs frequently generate common Javanese names such as Ayu, Dwi, Bayu, Eko, and Sari. For Sundanese, commonly produced names include Lia, Budi, Dewi, and Rina. While these names are widely used across Indonesia, the variation in honorific terms is limited across all models, only Pak and Bu, along with their formal variants Bapak and Ibu, appear consistently. This suggests that, although some models demonstrate surface-level lexical diversity, deeper sociolinguistic features such as honorific variation remain underrepresented.
@@ 5.1 Classification
5.1.1 Setup
We adopt the classification accuracy metric, as proposed in both Mostafazadeh et al. (2016) and Koto et al. (2022), defined as the ratio of correctly predicted instances to the total number of test cases. For our experiments, we fine-tune several models: Qwen 2.5 7B (Qwen et al., 2025), Llama 3.1 8B (Grattafiori et al., 2024), Gemma 2 9B (Rivière et al., 2024), SahabatAI Llama 8B (GoToCompany, 2024), SEA-LION Llama 8B (Ng et al., 2025), and XLM-R (Conneau et al., 2019). For LLMs, we conduct instruction fine-tuning with a multiple-choice question framework using a LoRA adapter Hu et al. (2022) on the combined Javanese and Sundanese training set. Full training details are provided in Appendix A. To ensure robustness, results are averaged over three runs. For comparison, we also report the zero-shot performance of each model prior to fine-tuning.
@@ 5.1.1 Setup
Table 3: Accuracy of models on the human-written test set in Javanese and Sundanese across training sets.
Model Javanese Sundanese
0-shot $_{\text{Train}} MT$$_{\text{Train}}$+GPT-4o All_LLM LLM_Filt 0-shot $_{\text{Train}} MT$$_{\text{Train}}$+GPT-4o All_LLM LLM_Filt
XLM-R NA 69.00 67.67 80.72 81.47 NA 75.42 50.34 75.93 72.73
Qwen 2.5 76.94 85.44 84.31 84.88 87.71 64.48 74.75 72.22 80.64 79.63
Llama 3.1 75.24 90.17 86.96 91.68 92.44 48.32 79.97 71.55 81.14 78.62
Gemma 2 85.07 93.95 95.46 95.46 95.27 51.52 86.20 89.23 87.88 85.02
SahabatAI Llama 77.13 95.46 94.14 97.16 97.73 34.51 87.88 87.21 91.25 92.42
SEA-LION Llama 69.19 96.60 94.71 93.76 95.27 25.93 88.55 83.67 82.15 77.44
Table 4: Accuracy of models on the machine-translated test set in Javanese and Sundanese across training sets.
Model Javanese Sundanese
0-shot $_{\text{Train}} MT$$_{\text{Train}}$+GPT-4o All_LLM LLM_Filt 0-shot $_{\text{Train}} MT$$_{\text{Train}}$+GPT-4o All_LLM LLM_Filt
XLM-R NA 69.40 70.20 66.80 62.60 NA 56.60 73.00 67.60 66.00
Qwen 2.5 68.00 80.00 75.00 75.80 75.80 63.60 78.80 74.20 76.60 77.60
Llama 3.1 47.60 86.00 70.40 77.40 73.40 48.60 83.40 74.00 75.00 76.00
Gemma 2 63.00 91.80 92.80 83.60 83.40 57.60 90.00 91.80 80.80 83.40
SahabatAI Llama 34.80 90.00 87.60 84.20 88.40 30.80 87.40 86.60 83.60 87.60
SEA-LION Llama 32.20 92.60 87.40 76.20 72.20 29.80 92.20 88.80 77.60 71.20
5.1.2 Overall Performance
Table 3 and Table 4 present the classification accuracies of all models trained on different training sets, evaluated on the human-written and machine-translated test sets, respectively. In the zero-shot setting, model performance ranges from approximately 34% to 70% across both Javanese and Sundanese, underscoring the inherent challenges of cultural commonsense reasoning in these languages. Fine-tuning consistently improves model performance over the zero-shot baseline, with LLMs showing particularly strong gains. Notably, XLM-R underperforms relative to the LLMs, suggesting that large generative models are more effective at capturing cultural nuances.
[…]
5.1.3 Performance Across Different Topics
[…]
@@ 5.2 Generation
5.2.1 Setup
Using the same set of LLMs from the classification experiments, we fine-tune the models to perform story continuation: given a four-sentence premise, the model is trained to generate a coherent fifth sentence in either Javanese or Sundanese. We apply supervised fine-tuning using QLoRA (Dettmers et al., 2023), with training details provided in Appendix A. To evaluate generation quality, we use two automatic metrics: ROUGE-L (Lin, 2004) for lexical overlap and BERTScore (Zhang et al., 2020) for semantic similarity.
5.2.2 Overall Performance
Table 5 and Table 6 presents the automated evaluation metrics for different training sets across Sundanese and Javanese on both human-written and machine-translated test sets. Finetuning consistently improves performance over the 0-shot. Notably, on the human-written data that contains more cultural nuanced story, model fine-tuned with LLM-generated data gives higher improvement compared to others.
Among the models, SEA-LION Llama achieves the highest score across most test, then followed closely by Gemma 2. The LLM_Filt data often matches or outperforms the All_LLM settings that uses more samples. Similar to classification, in human-written test set, LLM-generated-training data tends to be better than in machine-translated test set.
Table 5: BERTScore / ROUGE-L F1 of models on the human-written test set in Javanese and Sundanese across training sets.
Model Javanese Sundanese
[…]
Qwen 2.5 72.5 / 21.8 71.7 / 19.2 72.0 / 19.5 72.4 / 21.7 71.9 / 19.5 69.2 / 15.4 70.3 / 15.6 70.3 / 16.5 70.5 / 16.7 69.2 / 14.6
Llama 3.1 62.4 / 8.3 72.7 / 21.5 72.4 / 21.1 72.7 / 22.9 72.7 / 22.2 59.4 / 5.6 70.0 / 16.1 70.3 / 16.23 70.6 / 17.9 68.9 / 15.5
Gemma 2 70.9 / 17.7 73.2 / 24.0 73.1 / 23.1 72.8 / 22.9 72.6 / 22.2 68.2 / 13.2 70.3 / 18.1 70.2 / 17.9 70.8 / 17.9 70.2 / 17.1
SahabatAI Llama 63.1 / 15.3 71.8 / 20.1 59.7 / 9.2 62.6 / 13.6 67.2 / 16.6 57.7 / 8.9 69.2 / 16.2 59.1 / 7.5 62.4 / 11.7 70.6 / 17.6
SEA-LION Llama 72.7 / 23.0 72.4 / 20.9 72.5 / 21.5 72.8 / 23.9 72.6 / 22.5 68.9 / 16.7 70.1 / 16.3 70.3 / 17.0 71.0 / 18.3 71.3 / 18.6
Table 6: BERTScore / ROUGE-L F1 of models on the machine-translated test set in Javanese and Sundanese across training sets.
Model Javanese Sundanese
0-shot $_{\text{Train}} MT$$_{\text{Train}}$+GPT-4o All_LLM LLM_Filt 0-shot $_{\text{Train}} MT$$_{\text{Train}}$+GPT-4o All_LLM LLM_Filt
Qwen 2.5 70.5 / 15.6 70.7 / 18.3 71.0 / 18.1 71.6 / 18.8 70.6 / 15.4 69.9 / 14.0 70.9 / 17.8 70.9 / 18.1 71.3 / 16.7 70.2 / 13.7
Llama 3.1 59.7 / 4.4 72.1 / 21.4 72.0 / 20.7 71.8 / 20.4 71.3 / 18.8 59.3 / 4.2 71.4 / 19.2 71.2 / 18.6 71.3 / 17.7 68.8 / 15.7
Gemma 2 69.7 / 14.5 72.3 / 22.5 72.1 / 21.5 71.9 / 20.9 71.4 / 19.7 69.4 / 14.1 72.4 / 21.3 72.1 / 20.7 71.9 / 19.0 70.4 / 17.3
SahabatAI Llama 59.2 / 9.3 72.1 / 22.2 60.2 / 8.9 62.5 / 12.4 64.6 / 12.8 58.1 / 7.9 71.5 / 20.5 59.9 / 8.0 62.7 / 10.9 70.4 / 16.6
SEA-LION Llama 69.7 / 16.9 72.4 / 21.3 72.0 / 20.5 71.9 / 20.7 71.6 / 19.6 69.7 / 15.4 72.1 / 19.5 71.5 / 18.0 71.5 / 18.0 71.1 / 18.5
Table 7: Human evaluation results for Javanese and Sundanese.
Language Model Coherence Fluency Correctness Cultural Rel.
Javanese Gemma 2 4.92 4.94 76 92
SEA-LION Llama 4.98 4.98 84 92
Sundanese Gemma 2 4.22 4.74 74 86
SEA-LION Llama 4.40 4.92 78 94
5.2.3 Human Evaluation
We conducted a human evaluation to compare models between SEA-LION Llama and Gemma 2 fine-tuned on All LLM-generated data, the two models that showed strong performance in both classification and generation tasks (see Table 345, and  6). Using the human-written test set, annotators were presented with a story premise and the generated ending sentence. They were asked to rate the outputs based on coherence, fluency, correctness, and cultural relevance, following the same guidelines as in the earlier manual evaluation. As shown in Table 7, both models perform well in Javanese and Sundanese, with SEA-LION Llama slightly outperforming Gemma 2 across most human evaluation criteria. However, this advantage is less evident when measured using automatic metrics.
@@ 7 Limitations
This study acknowledges several limitations in terms of cultural nuance, language scope, and dialectal representation. While we carefully curated data across 12 cultural topics, our focus was limited to only two local languages—Javanese and Sundanese. Although these are the most widely spoken regional languages in Indonesia, they do not capture the full linguistic and cultural diversity present in the world. Moreover, our predefined topics and data sources may not comprehensively reflect the rich variation of cultural practices, dialects, and regional expressions within these languages.
@@ Appendix A Training Configurations
For classification, we set the maximum token length for the pre-trained language model to 450 for the premise and 50 for the ending sentence. The model was trained over 20 epochs with early stopping (patience set to 5), using a batch size of 40, Adam optimizer, an initial learning rate of 5e-6 for XLM-R, and a warm-up phase comprising 10% of the total training steps.
For training the text generation models, we use 4-bit quantization with a LoRA rank of 64 and a LoRA alpha of 128. The models are trained with a batch size of 8, gradient accumulation of 8, a learning rate of 2e-4, and for a single epoch. We employ the Unsloth.ai framework for efficient fine-tuning (Daniel Han and team, 2023).
For training the classifier for data filtering, we fine-tune the XLM-R model with a maximum token length of 1024 for the premise and 128 for the ending sentence. The model was trained over 26 epochs, using a batch size of 16, Adam optimizer, an initial learning rate of 1e-5 for XLM-R, and a warm-up phase comprising 5% of the total training steps.
@@
Appendix B Distribution of Filtered LLM-Generated Training Data
Initially, training data was generated using six different LLMs, with each model contributing approximately 16.67% of the total 10K samples (around 2K samples per model). However, after filtering the bad examples, the final dataset composition shifted. The data consist of 1,220 samples, the distribution is as follows: Claude (37.7%), GPT-4o (29.0%), LLama (8.6%), Cohere (4.3%), and Gemma-2 (1.7%).
Figure 4: Prompt template instructing LLM to generate a new example for Javanese and Sundanese.
Appendix C Agreement Score
We measure the fluency and coherency agreement scores using Pearson’s correlation and computed the correctness and cultural relevance scores using Cohen’s kappa. They are summarized in Table 8.
Table 8: Agreement Scores for Javanese and Sundanese